# RG-Invariant Symmetry Ratio for QCD: A Study of $U(1)_{A}$ and Chiral Symmetry Restoration

### v1

@@ 1 Introduction
The pattern of symmetry realization is a defining feature of any quantum field theory, governing its phase structure and the spectrum of its excitations. In Quantum Chromodynamics (QCD), the theory of the strong interaction, two global symmetries play a pivotal role: the chiral $SU(2)_{L}\times SU(2)_{R}$ symmetry of the light ($u$, $d$) quark sector, and the $U(1)_{A}$ axial symmetry. In the vacuum, the former is spontaneously broken by the chiral condensate $\langle\bar{q}q\rangle$ Nambu:1961tp ; Nambu:1961fr , giving mass to nucleons and giving rise to pions as pseudo-Goldstone bosons. The latter is explicitly broken by the axial anomaly, contributing significantly to the mass of the $\eta^{\prime}$ meson tHooft:1976snw ; Witten:1979vv ; Veneziano:1979ec . A cornerstone of modern nuclear physics is understanding how these broken symmetries behave under extreme conditions of temperature and density, such as those realized in heavy-ion collisions or within neutron stars.
@@
A profound and long-standing question is whether the effective restoration of the $U(1)_{A}$ symmetry, linked to the suppression of topological gauge fluctuations, coincides with this chiral crossover or occurs at a distinctly higher temperature $T_{1}>T_{c}$ Pisarski:1983ms ; Leutwyler:1992yt ; Cohen:1996ng ; Aoki:2012yj . Resolving this hierarchy is essential for a complete understanding of the QGP’s structure, the nature of the QCD transition, and the validity of effective models. Progress on this question via first-principles lattice QCD simulations has been challenging. A consistent picture among them has not emerged yet, see e.g., refs. Cossu:2013uua ; Buchoff:2013nra ; Brandt:2016daq ; Tomiya:2016jwr ; Ding:2020xlj ; Kaczmarek:2023bxb ; Aoki:2020noz ; Chiu:2023hnm ; Gavai:2024mcj , and the recent review by H.T. Ding at Lattice 2025, 2–8 November, Mumbai, India.
@@
Second, we present the first application of $\kappa_{AB}$ to the problem of symmetry restoration in QCD. We perform lattice simulations with $N_{f}=2+1+1$ optimal domain-wall quarks at the physical point, using three lattice spacings and twelve temperatures in the range 164–385 MeV. We compute $\kappa_{AB}$ across three distinct symmetry-breaking channels in the nonsinglet sector: the $U(1)_{A}$-breaking (scalar, pseudoscalar) channel ($\kappa_{PS}$) probing the $\pi$-$\delta$ system; the $SU(2)_{L}\times SU(2)_{R}$-breaking (vector, axial-vector) channel ($\kappa_{VA}$) probing the $\rho$-$a_{1}$ system; and an additional $U(1)_{A}$-sensitive channel using tensor–axial-tensor operators ($\kappa_{TX}$) which probes the $\rho$-$b_{1}$ system through the $\bar{q}\gamma_{4}\gamma_{k}q$ and $\bar{q}\gamma_{5}\gamma_{4}\gamma_{k}q$ currents. The inclusion of the tensor channel provides a vital cross-check, as it probes the $U(1)_{A}$ anomaly through a different Dirac structure than the scalar channel.
Our key findings are as follows. At finite lattice spacing, we observe a clear ordering: the breaking in the $U(1)_{A}$ scalar channel is strongest, while the $U(1)_{A}$ tensor channel and the $SU(2)$ vector channel are significantly weaker and comparable to each other ($\kappa_{PS}>\kappa_{TX}\sim\kappa_{VA}$). However, a controlled continuum extrapolation reveals that this entire hierarchy collapses. All three symmetry-breaking strengths become statistically indistinguishable within our resolution. This result provides a new, high-precision benchmark from a chirally symmetric action. It indicates that the effective restoration scales probed by these different operators converge in the continuum limit, suggesting that the manifestations of $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ symmetries in the nonsinglet sector are much more concurrent than suggested by finite-lattice-spacing studies focused on a single channel.
The paper is organized as follows. In Sec. 2, we formally define the RG-invariant symmetry ratio $\kappa_{AB}$ and detail its theoretical properties. Section 3 describes our lattice setup, time-correlation functions, and analysis methodology. Our numerical results at finite lattice spacing are presented in Sec. 4, and the crucial continuum extrapolation is performed in Sec. 5. In Sec. 6, we discuss the implications of our findings, and outline future applications of the $\kappa_{AB}$ framework.
@@ 2.1 Correlation functions and symmetry partners
where $\Gamma$ is a Dirac matrix specifying the quantum numbers of the channel, $t^{a}$ (with $a=1,\dots,N_{f}^{2}-1$) are the generators of $SU(N_{f})$ for flavor nonsinglets, and $t^{0}\equiv\mbox{1\hskip-3.41432ptI}=\text{diag}(1,1,\dots,1)/\sqrt{2N_{f}}$ (identity operator) for the flavor singlet, with the normalization condition ${\rm tr}(t^{a}t^{b})=\delta^{ab}/2$ ($a=0,\dots,N_{f}^{2}-1$). For flavor-singlet operators we adopt the shorthand notation $\bar{q}\Gamma q\equiv\bar{q}\Gamma t^{0}q$. The factor $t^{0}$ is suppressed in the notation but always implied; this convention is used throughout this work. We use the notation $(x_{1},x_{2},x_{3},x_{4})\equiv(x,y,z,t)\equiv(\vec{x},t)$ interchangeably. The $t$-correlator is defined as
@@
Composite operators such as $O_{\Gamma}^{a}(x)$ require regularization and renormalization. Both $C_{\Gamma}(t)$ and $C_{\Gamma}(z)$ contain short-distance singularities; for instance, they diverge as $t^{-3}$ or $z^{-3}$ when the two operators coincide. In what follows we concentrate on the $t$-correlator; the extension to the $z$-correlator is straightforward.
@@ 2.2 Renormalization and RG invariance
  • under $U(1)_{A}$ (tensor channel): the tensor vector ($T$, e.g. $\rho^{+}\sim\bar{d}\gamma_{4}\gamma_{i}u$) and axial-tensor vector ($X$, e.g. $b_{1}^{+}\sim\bar{d}\gamma_{5}\gamma_{4}\gamma_{i}u$) are partners, hence $Z_{T}=Z_{X}$.
@@ 2.3 Integrated spectral weights and the symmetry strength parameter
Although $\kappa_{AB}(t)$ is a valid probe, its dependence on $t$ and the statistical fluctuations at individual times make a channel-wide comparison of symmetry breaking cumbersome. A more robust global measure is obtained by integrating over Euclidean time, which sums the spectral weight in each channel. We therefore define the regularized susceptibility for channel $A$ by subtracting the singular contact term at $t=0$, and for the flavor singlet and the $N_{f}-1$ flavor nonsinglets with Cartan (diagonal) generators that contain disconnected diagrams, also the vacuum contribution.
$\chi_{A}=\int_{t\neq 0}dt\left[C_{A}(t)-\int d^{3}x\langle\bar{q}\Gamma_{A}t^{a}q\rangle_{\text{vev}}^{2}\right],$ (4)
@@
where $\chi_{A}\geq\chi_{B}>0$, and $0\leq\kappa_{AB}\leq 1$. This parameter provides an intuitive scale: $\kappa_{AB}=0$ indicates exact degeneracy of the integrated spectral weights (the symmetry is manifest), whereas $\kappa_{AB}\to 1$ signals maximal asymmetry. The construction is general and can be applied to study any symmetry in quantum field theory for which partner operators can be defined.
Remark on terminology. Throughout this work, “restoration” of a symmetry always refers to effective restoration — i.e., the corresponding RG‑invariant ratio $\kappa_{AB}$ becomes consistent with zero within statistical uncertainties. Exact restoration, which would require both vanishing quark masses and the complete absence of the $U(1)_{A}$ anomaly, is neither achievable in physical QCD nor required for the validity of our conclusions. This operational definition is implicit in all subsequent uses of “degeneracy”, restoration”, and $\kappa_{AB}=0$”.
@@ 2.4 Symmetry Channels Involving Flavor Singlet Operators
For $N_{f}=2$ QCD, effective restoration of $SU(2)_{L}\times SU(2)_{R}$ chiral symmetry implies equalities among regularized susceptibilities as shown in eqs. (7)–(11),
@@
$\chi^{s}_{T_{k}}$ $=\chi^{ns}_{X_{k}}\Longleftrightarrow\chi_{\omega}=\chi_{b_{1}},\qquad k=1,2,3,$ (9)
$\chi^{s}_{X_{k}}$ $=\chi^{ns}_{T_{k}}\Longleftrightarrow\chi_{h_{1}}=\chi_{\rho},\qquad k=1,2,3,$ (10)
@@
where the superscripts $s$ and $ns$ denote singlet and nonsinglet channels, respectively. The notations of meson operators and their corresponding regularized susceptibilies are summarized in appendix A
@@
$\chi^{ns}_{T_{k}}$ $=\chi^{ns}_{X_{k}}\Longleftrightarrow\chi_{\rho}=\chi_{b_{1}},\qquad k=1,2,3.$ (14)
$\chi^{s}_{T_{k}}$ $=\chi^{s}_{X_{k}}\Longleftrightarrow\chi_{\omega}=\chi_{h_{1}},\qquad k=1,2,3,$ (15)
@@
In particular, under $SU(2)_{L}\times SU(2)_{R}$ chiral rotations, the operators transform as
  • $(\sigma,\pi^{a})$: singlet scalar $\sigma\equiv\bar{q}q$ with nonsinglet pseudoscalar $\pi^{a}\equiv\bar{q}\gamma_{5}t^{a}q$,
  • $(\eta,\delta^{a})$: singlet pseudoscalar $\eta\equiv\bar{q}\gamma_{5}q$ with nonsinglet scalar $\delta^{a}\equiv\bar{q}t^{a}q$.
In principle, each pair could be probed by the symmetry ratio $\kappa_{AB}$ defined in Eq. 6, offering additional insight into chiral restoration.
In a lattice formulation with exact chiral symmetry, the transformation laws enforce relationships between the renormalization constants of these operators, e.g., $Z_{S}^{s}=Z_{P}^{ns}$ for $(\sigma,\pi)$ and $Z_{P}^{s}=Z_{S}^{ns}$ for $(\eta^{\prime},\delta)$, where $s$ and $ns$ denote singlet and nonsinglet, respectively. Despite these formal relations, several practical obstacles prevent a clean extraction of $\kappa_{AB}$ for singlet-involved channels in the present study:
@@
    Correlators of the singlet operators $\sigma$ and $\eta^{\prime}$ receive contributions from disconnected Wick contractions, which are notoriously noisy in lattice simulations and require extensive stochastic sampling. By contrast, their nonsinglet partners involve only connected diagrams. This asymmetry in diagrammatic structure leads to very different signal-to-noise ratios and systematic uncertainties in the estimation of the susceptibilities $\chi_{A}$ and $\chi_{B}$, compromising a direct comparison.
@@
    Power divergences and vacuum subtraction.
    The singlet scalar operator $\bar{q}q$ acquires a power-divergent vacuum expectation value (vev) $\langle\bar{q}q\rangle\neq 0$, requiring a vacuum subtraction in the renormalized susceptibility:
    $\chi_{\sigma}=\sum_{t\neq 0}\sum_{\vec{x}}\bigl[\langle\bar{q}q(\vec{x},t)\,\bar{q}q(0)\rangle-\langle\bar{q}q\rangle^{2}\bigr].$
    For the singlet pseudoscalar $\bar{q}\gamma_{5}q$, while parity symmetry ensures $\langle\bar{q}\gamma_{5}q\rangle=0$ in the continuum, its lattice renormalization mixes with the topological charge density $F\tilde{F}$ through power-divergent coefficients $\sim 1/a^{n}$. This induces lattice artifacts that can mimic a nonzero vacuum expectation, and a careful subtraction is needed to isolate the physical signal. These subtractions introduce systematic uncertainties that propagate into symmetry ratios such as $\kappa_{\sigma-\pi}$ and $\kappa_{\eta^{\prime}-\delta}$, whereas nonsinglet partners avoid such complications.
@@
    Anomalous renormalization of susceptibilities of singlet scalar and pseudoscalar mesons
    The regularized susceptibilities of flavor-singlet scalar ($\sigma\sim\bar{q}q$) and pseudoscalar ($\eta^{\prime}\sim\bar{q}\gamma_{5}q$) mesons inherit anomalous renormalization patterns due to the $U(1)_{A}$ anomaly Adler:1969gk ; Bell:1969ts ; Fujikawa:1979ay .
    In this subsection, all dimensionful quantities are expressed in lattice units with $a=1$. For the scalar singlet channel, the renormalized susceptibility $\chi_{\sigma}^{R}$ satisfies
    $\chi_{\sigma}^{R}=(Z_{S}^{s})^{2}\chi_{\sigma}^{\text{bare}}+b_{S}\,m_{q}+c_{S}\,\chi_{t}^{\text{bare}}+\text{finite mixings},$ (16)
    where $b_{S}$ and $c_{S}$ are dimensionless, regularization- and scheme-dependent coefficients Martinelli:1995ty . Here $\chi_{t}^{\text{bare}}\equiv\langle Q_{t}^{2}\rangle/V$ is the bare topological susceptibility in lattice units. The term $\propto m_{q}$ encodes power-divergent mixing with the identity, while the term $\propto\chi_{t}^{\text{bare}}$ embodies power-divergent mixing with the topological susceptibility.
    Importantly, for overlap fermions with exact chiral symmetry, all power-divergent mixings vanish identically; only finite, anomaly-induced mixing survives in the continuum limit. For domain‑wall fermions with finite $N_{s}$, power divergences are exponentially suppressed $\sim e^{-\alpha N_{s}}$ and become negligible in the continuum extrapolation. After power divergences are subtracted, the surviving finite part of $c_{S}$ becomes a scale-independent, finite constant in the continuum limit.
    For the pseudoscalar singlet susceptibility $\chi_{\eta^{\prime}}^{R}$:
    $\chi_{\eta^{\prime}}^{R}=(Z_{P}^{s})^{2}\,\chi_{\eta^{\prime}}^{\text{bare}}+\tilde{c}(\mu)\,\frac{\alpha_{s}}{4\pi}\,\chi_{t}^{R}+\mathcal{O}(a),$ (17)
    where $Z_{P}^{s}=Z_{S}^{ns}$ for chirally symmetric fermions. Here $\tilde{c}(\mu)$ is a dimensionless but scale-dependent mixing coefficient required by the axial anomaly. Its renormalization group running
    $\tilde{c}(\mu)=\tilde{c}(\mu_{0})\left(\frac{\alpha_{s}(\mu)}{\alpha_{s}(\mu_{0})}\right)^{\gamma_{0}/\beta_{0}},$
    follows from the anomalous dimension $\gamma_{0}$ of the singlet pseudoscalar operator. In domain‑wall fermion actions, residual chiral symmetry breaking can induce a mild $\mathcal{O}(a)$ mixing with the identity, but this is parametrically smaller than the anomaly-induced finite mixing with $\chi_{t}$.
    Constructing fully RG-invariant symmetry ratios such as $\kappa_{\sigma-\pi}$, $\kappa_{\eta-\delta}$, or $\kappa_{\sigma-\eta}$ would therefore require a consistent non‑perturbative renormalization that disentangles the power-divergent and anomaly-induced mixings, coupled with high-precision measurements of noisy disconnected diagrams—a substantial task that we defer to future work.

### v2

@@ 1 Introduction
The pattern of symmetry realization is a defining feature of any quantum field theory, governing its phase structure and the spectrum of its excitations. In Quantum Chromodynamics (QCD), the theory of the strong interaction, two global symmetries play a pivotal role: the chiral $SU(2)_{L}\times SU(2)_{R}$ symmetry of the light ($u$, $d$) quark sector, and the $U(1)_{A}$ axial symmetry. In the vacuum, the former is spontaneously broken by the chiral condensate $\langle\bar{q}q\rangle$ Nambu:1961tp ; Nambu:1961fr , giving mass to nucleons and generating pions as pseudo-Goldstone bosons. The latter is explicitly broken by the axial anomaly Adler:1969gk ; Bell:1969ts ; Fujikawa:1979ay , contributing significantly to the mass of the $\eta^{\prime}$ meson tHooft:1976snw ; Witten:1979vv ; Veneziano:1979ec . A cornerstone of modern nuclear physics is understanding how these broken symmetries behave under extreme conditions of temperature and density, such as those realized in heavy-ion collisions or within neutron stars.
@@
A profound and long-standing question is whether the effective restoration of the $U(1)_{A}$ symmetry, linked to the suppression of topological gauge fluctuations, coincides with this chiral crossover or occurs at a distinctly higher temperature $T_{1}>T_{c}$ Pisarski:1983ms ; Leutwyler:1992yt ; Cohen:1996ng ; Aoki:2012yj . Resolving this hierarchy is essential for a complete understanding of the QGP’s structure, the nature of the QCD transition, and the validity of effective models. Progress on this question via first-principles lattice QCD simulations has been challenging. A consistent picture among them has not emerged yet, see e.g., refs. Cossu:2013uua ; Buchoff:2013nra ; Brandt:2016daq ; Tomiya:2016jwr ; Ding:2020xlj ; Kaczmarek:2023bxb ; Aoki:2020noz ; Chiu:2023hnm ; Gavai:2024mcj , and the recent review by H.T. Ding at Lattice 2025222https://indico.global/event/14504/contributions/137953/, 2–8 November, Mumbai, India.
@@
Second, we present the first application of $\kappa_{AB}$ to the problem of symmetry restoration in QCD. We perform lattice simulations with $N_{f}=2+1+1$ optimal domain-wall quarks at the physical point, using three lattice spacings and twelve temperatures in the range 164–385 MeV. We compute $\kappa_{AB}$ across three distinct symmetry-breaking channels in the nonsinglet sector: the $U(1)_{A}$-breaking (scalar, pseudoscalar) channel ($\kappa_{PS}$) probing the $\pi$-$\delta$ system; the $SU(2)_{L}\times SU(2)_{R}$-breaking (vector, axial-vector) channel ($\kappa_{VA}$) probing the $\rho$-$a_{1}$ system; and an additional $U(1)_{A}$-sensitive channel using tensor–axial-tensor operators ($\kappa_{TX}$) which probes the $\rho_{T}$-$b_{1}$ system through the $\bar{q}\gamma_{4}\gamma_{k}q$ and $\bar{q}\gamma_{5}\gamma_{4}\gamma_{k}q$ currents. The inclusion of the tensor channel provides a vital cross-check, as it probes the $U(1)_{A}$ anomaly through a different Dirac structure than the scalar channel.
Our key findings are as follows. At finite lattice spacing, we observe a clear ordering: the breaking in the $U(1)_{A}$ scalar–pseudoscalar channel is strongest, while the $U(1)_{A}$ tensor–axial-tensor channel and the $SU(2)_{L}\times SU(2)_{R}$ vector–axial-vector channel are significantly weaker and comparable to each other ($\kappa_{PS}>\kappa_{TX}\sim\kappa_{VA}$). However, a controlled continuum extrapolation reveals that this entire hierarchy collapses. All three symmetry-breaking strengths become statistically indistinguishable within our resolution. This result provides a new, high-precision benchmark from a chirally symmetric action. It indicates that the effective restoration scales probed by these different operators converge in the continuum limit, suggesting that the manifestations of $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ symmetries in the nonsinglet sector are much more concurrent than suggested by finite-lattice-spacing studies focused on a single channel.
The paper is organized as follows. In section 2, we formally define the RG-invariant symmetry ratio $\kappa_{AB}$ and detail its theoretical properties. Our lattice setup, time-correlation functions, and analysis methodology are described in section 3. Numerical results at finite lattice spacing are presented in section 4, and the continuum extrapolation is performed in section 5. The implications of our findings and future applications of the $\kappa_{AB}$ framework are discussed in section 6.
@@ 2.1 Correlation functions and symmetry partners
where $\Gamma$ is a Dirac matrix specifying the quantum numbers of the channel, $t^{a}$ (with $a=1,\dots,N_{f}^{2}-1$) are the generators of $SU(N_{f})$ for flavor nonsinglets, and $t^{0}\equiv\mbox{1\hskip-3.41432ptI}=\text{diag}(1,1,\dots,1)/\sqrt{2N_{f}}$ (identity operator) for the flavor singlet, with the normalization condition ${\rm tr}(t^{a}t^{b})=\delta^{ab}/2$ ($a=0,\dots,N_{f}^{2}-1$). For flavor-singlet operators we adopt the shorthand notation $\bar{q}\Gamma q\equiv\bar{q}\Gamma t^{0}q$, with the factor $t^{0}$ always implied but suppressed. We use the notation $(x_{1},x_{2},x_{3},x_{4})\equiv(x,y,z,t)\equiv(\vec{x},t)$ interchangeably. With this notation, the $t$-correlator is defined as
@@
Composite operators such as $O_{\Gamma}^{a}(x)$ require regularization and renormalization. Both $C_{\Gamma}(t)$ and $C_{\Gamma}(z)$ contain short-distance singularities; the leading divergence behaves as $t^{-3}$ or $z^{-3}$ in the continuum limit as the two operators approach coincidence, corresponding to a contact term $\sim a^{-3}$ on the lattice whose precise form is regularization-dependent. In what follows we concentrate on the $t$-correlator; the extension to the $z$-correlator is straightforward.
@@ 2.2 Renormalization and RG invariance
  • under $U(1)_{A}$ (tensor channel): the tensor ($T$, e.g. $\rho_{T}^{+}\sim\bar{d}\gamma_{4}\gamma_{i}u$) and axial-tensor ($X$, e.g. $b_{1}^{+}\sim\bar{d}\gamma_{5}\gamma_{4}\gamma_{i}u$) are partners, hence $Z_{T}=Z_{X}$.
@@ 2.3 Integrated spectral weights and the symmetry strength parameter
Although $\kappa_{AB}(t)$ is a valid probe, its dependence on $t$ and the statistical fluctuations at individual times make a channel-wide comparison of symmetry breaking cumbersome. A more robust global measure is obtained by integrating over Euclidean time, which sums the spectral weight in each channel. We therefore define the regularized susceptibility for channel $A$. The contact term at $t=0$ is subtracted to remove the short-distance singularity; for flavor-singlet and Cartan nonsinglet channels with scalar or tensor Dirac structure (for which the vacuum expectation value $\langle\bar{q}\Gamma_{A}t^{a}q\rangle$ is nonzero), the vacuum contribution is also subtracted. The integration runs over $t\in(0,1/T)$, where $T$ is the temperature, with the $t=0$ contact term excluded:
$\chi_{A}=\int_{0<t<1/T}dt\left[C_{A}(t)-\int d^{3}x\langle\bar{q}\Gamma_{A}t^{a}q\rangle_{\text{vev}}^{2}\right],$ (4)
@@
where $\chi_{A}\geq\chi_{B}>0$ (with the convention $\chi_{A}\geq\chi_{B}$ adopted without loss of generality), and $0\leq\kappa_{AB}\leq 1$. This parameter provides an intuitive scale: $\kappa_{AB}=0$ indicates exact degeneracy of the integrated spectral weights (the symmetry is manifest), whereas $\kappa_{AB}\to 1$ signals maximal asymmetry. The construction is general and can be applied to study any symmetry in quantum field theory for which partner operators can be defined.
Remark on terminology. Throughout this work, “restoration” of a symmetry always refers to effective restoration — i.e., the corresponding RG‑invariant ratio $\kappa_{AB}$ becomes consistent with zero within statistical uncertainties. Exact restoration, which would require vanishing quark masses for $SU(2)_{L}\times SU(2)_{R}$ and additionally the complete suppression of the $U(1)_{A}$ anomaly for $U(1)_{A}$ restoration, is neither achievable in physical QCD nor required for the validity of our conclusions. This operational definition is implicit in all subsequent uses of “degeneracy”, “restoration”, and $\kappa_{AB}=0$”.
@@ 2.4 Symmetry Channels Involving Flavor Singlet Operators
For $N_{f}=2$ massless QCD, effective restoration of $SU(2)_{L}\times SU(2)_{R}$ chiral symmetry implies equalities among regularized susceptibilities as shown in eqs. (7)–(11),
@@
$\chi^{s}_{T_{k}}$ $=\chi^{ns}_{X_{k}}\Longleftrightarrow\chi_{\omega_{T}}=\chi_{b_{1}},\qquad k=1,2,3,$ (9)
$\chi^{s}_{X_{k}}$ $=\chi^{ns}_{T_{k}}\Longleftrightarrow\chi_{h_{1T}}=\chi_{\rho_{T}},\qquad k=1,2,3,$ (10)
@@
where the superscripts $s$ and $ns$ denote singlet and nonsinglet channels, respectively. Recall that the physical $\eta^{\prime}$ corresponds to the $\eta$ singlet in $N_{f}=2$ QCD. The notations of meson operators and their corresponding regularized susceptibilies are summarized in appendix A.
@@
$\chi^{ns}_{T_{k}}$ $=\chi^{ns}_{X_{k}}\Longleftrightarrow\chi_{\rho_{T}}=\chi_{b_{1}},\qquad k=1,2,3.$ (14)
$\chi^{s}_{T_{k}}$ $=\chi^{s}_{X_{k}}\Longleftrightarrow\chi_{\omega_{T}}=\chi_{h_{1T}},\qquad k=1,2,3,$ (15)
@@
Equality of renormalization constants from symmetry. For chirally symmetric fermion formulations, the symmetry relations in eqs. (7)–(15) imply specific equalities among renormalization constants.
For scalar and pseudoscalar channels, the partner relations yield:
  • From $SU(2)_{L}\times SU(2)_{R}$: $(\sigma,\pi)$ and $(\eta,\delta)$ are partners, giving $Z_{S}^{s}=Z_{P}^{ns}$ and $Z_{P}^{s}=Z_{S}^{ns}$.
  • From $U(1)_{A}$: $(\delta,\pi)$ and $(\sigma,\eta)$ are partners, giving $Z_{S}^{ns}=Z_{P}^{ns}$ and $Z_{S}^{s}=Z_{P}^{s}$.
Combining these gives:
$Z_{S}^{ns}=Z_{P}^{ns}=Z_{S}^{s}=Z_{P}^{s}.$
For tensor and axial-tensor channels, the same logic applies:
  • Under $SU(2)_{L}\times SU(2)_{R}$: $(T^{s},X^{ns})$ and $(X^{s},T^{ns})$ are partners, giving $Z_{T}^{s}=Z_{X}^{ns}$ and $Z_{X}^{s}=Z_{T}^{ns}$.
  • Under $U(1)_{A}$: $(T^{s},X^{s})$ and $(T^{ns},X^{ns})$ are partners, giving $Z_{T}^{s}=Z_{X}^{s}$ and $Z_{T}^{ns}=Z_{X}^{ns}$.
Combining these gives:
$Z_{T}^{ns}=Z_{X}^{ns}=Z_{T}^{s}=Z_{X}^{s}.$
For vector and axial-vector channels, the pattern is different:
  • Under $SU(2)_{L}\times SU(2)_{R}$, the nonsinglet vector $\rho$ and axial-vector $a_{1}$ are partners, giving $Z_{V}^{ns}=Z_{A}^{ns}$. Both currents are conserved for all $m_{q}$, so $Z_{V}^{ns}=Z_{A}^{ns}=1$.
  • The singlet vector $\omega\sim\bar{q}\gamma_{\mu}q$ is associated with the exact $U(1)_{V}$ symmetry of the QCD action, which is preserved for all $m_{q}$, so $Z_{V}^{s}=1$.
  • The singlet axial-vector $h_{1\mu}\sim\bar{q}\gamma_{5}\gamma_{\mu}q$ is not conserved due to the $U(1)_{A}$ anomaly, regardless of quark mass. Its divergence is $\partial^{\mu}h_{1\mu}=2m_{q}\bar{q}\gamma_{5}q+\frac{g^{2}}{16\pi^{2}}F\tilde{F}$. Unlike the conserved currents above, $h_{1\mu}$ is not protected by a conserved charge, so its renormalization constant $Z_{A}^{s}$ is not fixed to unity and must be determined non-perturbatively. The current itself renormalizes multiplicatively (without operator mixing), but $Z_{A}^{s}\neq Z_{V}^{s}$ and $Z_{A}^{s}\neq 1$ in general.
These relations ensure that, after appropriate contact term and vacuum subtractions, any ratio $\kappa_{AB}$ involving operators from the same renormalization constant multiplet is RG-invariant. This includes the nonsinglet ratios $\kappa_{PS}$, $\kappa_{VA}$, $\kappa_{TX}$ studied in this work, as well as singlet-involved ratios such as $\kappa_{\sigma,\pi}$, $\kappa_{\eta,\delta}$, $\kappa_{\omega_{T},b_{1}}$, $\kappa_{h_{1T},\rho_{T}}$, and $\kappa_{\omega_{T},h_{1T}}$. Consequently, ratios like $\kappa_{\sigma,\pi}$ and $\kappa_{\eta,\delta}$ are formally RG-invariant, just like their nonsinglet counterparts.
Despite this formal RG invariance, several practical obstacles prevent a clean extraction of $\kappa_{AB}$ for channels involving flavor-singlet operators in the present study:
@@
    Correlators of the singlet operators $\sigma$ and $\eta^{\prime}$ receive contributions from disconnected Wick contractions, which are computationally demanding and statistically noisy in lattice simulations and require extensive stochastic sampling. By contrast, their nonsinglet partners involve only connected diagrams. This asymmetry in diagrammatic structure leads to very different signal-to-noise ratios and systematic uncertainties in the estimation of the susceptibilities $\chi_{A}$ and $\chi_{B}$, compromising a direct comparison.
@@
    Regularization of scalar and pseudoscalar singlet susceptibilities.
    Following eq. (4), the regularized susceptibilities for singlet channels are defined by subtracting the power-divergent contact term. For the scalar singlet $\sigma\sim\bar{q}q$, this requires subtracting the vacuum expectation value $\langle\bar{q}q\rangle^{2}$ to remove the power divergence $\propto 1/a^{2}$. For the pseudoscalar singlet $\eta^{\prime}\sim\bar{q}\gamma_{5}q$, parity ensures $\langle\bar{q}\gamma_{5}q\rangle=0$ in the continuum, so no vacuum subtraction is needed. However, the unsubtracted bare susceptibility contains power-divergent mixing with the topological charge density $F\tilde{F}$ through coefficients $\sim 1/a^{n}$. The contact term subtraction in eq. (4) removes these power-divergent mixings as well. After this subtraction, the resulting regularized susceptibility $\chi_{\eta^{\prime}}^{\mathrm{sub}}$ is free of all power divergences and multiplicatively renormalizable with $Z_{P}^{s}=Z_{S}^{ns}$. Crucially, all power-divergent mixings, including any power-divergent mixing with the topological susceptibility $\chi_{t}$, are completely removed by these subtractions.
@@
    Finite mixing with topological susceptibility.
    Even after all power divergences are removed, the $U(1)_{A}$ anomaly induces finite physical mixing between the singlet susceptibilities and the topological susceptibility $\chi_{t}$. For the scalar singlet, this mixing occurs through a scale-independent coefficient:
    $\chi_{\sigma}^{R}=Z_{S}^{2}\chi_{\sigma}^{\mathrm{sub}}+c_{S}\,\chi_{t}^{R},$
    where $c_{S}$ is a finite, scheme-dependent constant; its scale independence follows from the fact that $\chi_{\sigma}$ and $\chi_{t}$ share the same $CP$ quantum numbers and their mixing is mediated by a dimension-four operator whose anomalous dimension vanishes at leading order. For the pseudoscalar singlet, the mixing is through a scale-dependent coefficient:
    $\chi_{\eta^{\prime}}^{R}=Z_{S}^{2}\chi_{\eta^{\prime}}^{\mathrm{sub}}+\tilde{c}(\mu)\,\chi_{t}^{R},$
    with $\tilde{c}(\mu)$ evolving under the renormalization group due to the interplay with the axial anomaly Ward identity. These mixing coefficients are finite — they are not power-divergent and cannot be removed by the subtractions of eq. (4). Obtaining the fully renormalized susceptibilities $\chi_{\sigma}^{R}$ and $\chi_{\eta^{\prime}}^{R}$ therefore requires a non-perturbative determination of $c_{S}$ and $\tilde{c}(\mu)$ Martinelli:1995ty .
    Constructing fully RG-invariant ratios such as $\kappa_{\sigma,\pi}$ and $\kappa_{\eta,\delta}$ would then require performing a non-perturbative subtraction of the finite mixing terms. We defer this substantial task to future work.
    For overlap fermions with exact chiral symmetry, the topological charge can be defined via the index theorem ($Q_{t}=n_{+}-n_{-}$) through zero mode counting. In this case, $\chi_{t}$ requires no renormalization and the bare susceptibility equals the renormalized one: $\chi_{t}^{\mathrm{bare}}=\chi_{t}^{R}$. This simplifies the analysis, but the challenge of determining the finite mixing coefficients remains. In practice, at finite lattice spacing, gauge field dislocations can produce spurious near-zero modes that contaminate the zero-mode count; sufficiently smooth configurations are required for the index theorem identification to be unambiguous.